Reports 1-1 of 1 Clear search Modify search
CAL (General)
dan.chen - 8:34 Wednesday 28 January 2026 (36236) Print this report
Calibration after coil balance adjustment

CAL group
We did the calibration measurements and updated the parameters.

Estimated parameters in the Pre-maintenance measurements are as follows.
H_etmxtm = 3.235833627e-14 @10Hz ( -0.42% from previous measurements)
H_etmxim = 9.563757544e-15 @10Hz ( -1.31% from previous measurements)
Optical_gain = 3.726146467e+12 ( 23.03% from previous measurements)
Cavity_pole = 16.877398235 Hz ( 1.64% from previous measurements)

This time, a modification to K1LSC-OMC_DC FB was applied to the calibration parameters.
In this FB, a gain of 0.8 had previously been input, but its effect was moved to FM3 yesterday.
In the reconstruction, the gain is included in the optical gain part, while the FMs are included in the filter part.
Therefore, in order to adapt to this FB change, the reconstruction filter was modified.
As a result, the optical gain value increased by ~2 dB compared to the previous (yesterday’s) value.
However, this operation does not change the combined response of the filter and optical gain (there is a slight difference between yesterday’s and today’s measurements).

Previous values are listed in klog#36224.
 

Images attached to this report
Comments to this report:
Shingo Hido - 18:34 Monday 02 February 2026 (36261) Print this report

Sorry for the confusion. this is probably my mistake. I likely forgot to load the calibration filter.

Based on what I remember, the unloaded part is most likely TM_LOCL_L part.
I’m sure I had confirmed the other filters were loaded, so TM_LOCL_L is the most plausible one that was missed.

Also, I think that the current parameters have already been updated using the results from klog#36236.

Search Help
×

Warning

×